# A quasi-linear model of electromagnetic turbulent transport and its application to flux-driven transport predictions for STEP

### v1

@@ 1 Introduction
High-fidelity tools to evolve plasma equilibrium profiles on the transport time scale have been developed in recent years by coupling transport solvers to local or global gyrokinetic codes. These codes, such as GENE-TANGO (Di Siena et al., 2022), TGYRO-CGYRO (Candy et al., 2009) and TRINITY (Barnes et al., 2010), exploit the enormous separation between the timescales associated with turbulence and equilibrium transport, to evolve plasma profiles self-consistently with the turbulence. Naturally this approach requires considerably more computational resources than using reduced transport models [^1] This is true even when Gaussian process regression techniques are used to reduce the number of solver iterations (Rodriguez-Fernandez et al., 2022).. The application of such a high-fidelity gyrokinetic transport solver to STEP is computationally prohibitively expensive, because each of the many nonlinear local simulations required by a transport solver costs more than $10^{5}$ core-hours. Furthermore, starting from an initial state that strongly violates power balance increases significantly the number of nonlinear gyrokinetic simulations required for profile convergence (power balance is very poorly satisfied at several radial locations in STEP-EC-HD as assumed in scenario design with JINTRAC, which provides the initial condition in our flux-driven transport calculations).
@@ 2.1 Linear simulation results
Figure 1: Growth rate (a) and mode frequency (b) as functions of the binormal wave vector $k_{y}$ at various radial surfaces of STEP-EC-HD.
@@
On the innermost surface at $\Psi_{n}=0.15$, however, the dominant modes have negative frequency and very different parallel mode structures. These are illustrated in Fig. 3 [(a) and (b)] for the fastest growing mode at $k_{y}\rho_{s}\simeq 0.84$: $\delta\phi$ has odd parity and is highly extended in $\theta$ whilst $\delta A_{\parallel}$ is very localised at $\theta=0$ with even parity. Amplitudes of $e\delta\phi/T_{e}$ and $\delta A_{\parallel}/(\rho_{s}B_{0})$ are comparable, indicating that the mode is strongly electromagnetic. These negative frequency modes propagate in the electron diamagnetic drift direction, and have the typical properties expected of MTMs (Applegate et al., 2004). On the same surface at the lowest $k_{y}$, on the other hand, the dominant mode switches to a positive frequency and very different parallel mode structures in $\delta\phi$ and $\delta A_{\parallel}$. The parallel mode structure of a low $k_{y}$ mode with positive frequency is shown in Fig. 3 [(c) and (d)]: the parallel structure closely resemble Fig. 2, indicating that the hybrid-KBM is linearly dominant at low $k_{y}$. Although these low $k_{y}$ hybrid-KBMs are only weakly unstable, they nevertheless contribute significantly to the turbulent transport, as will be discussed in Sec. 2.2.
@@ 3.1 Quasi-linear reduced transport model
Species heat and particle fluxes are obtained from the total model heat flux $Q_{\mathrm{ql}}=Q_{0}\Lambda^{\alpha}$ defined in Eq. (2), using quasi-linear weights $Q_{\mathrm{l},s}/Q_{\mathrm{l}}$ and $\Gamma_{\mathrm{l},s}/Q_{\mathrm{l}}$, where $Q_{\mathrm{l},s}$ and $\Gamma_{\mathrm{l},s}$ are the linear heat and particle fluxes for species $s$, and $Q_{\mathrm{l}}=\sum_{s}Q_{\mathrm{l},s}$ is the total linear heat flux. Quasi-linear weights depend on $(k_{y},\theta_{0})$, especially when the character of the dominant instability changes with $k_{y}$ and $\theta_{0}$, therefore they are included in Eq. (5) before integrating over $k_{y}$ and $\theta_{0}$ to obtain species heat and particle fluxes:
@@ 3.2 Application to STEP relevant regimes
The parameters $Q_{0}$ and $\alpha$, which link $\Lambda$ to the quasi-linear heat flux $Q_{\mathrm{ql}}$ via Eq. (2), are tuned here to optimise the description of the total heat transport from hybrid-KBM turbulence in STEP. Suitable values are determined by fitting the nonlinear saturated heat fluxes from a set of nonlinear gyrokinetic simulations for STEP relevant local equilibria. This database contains nonlinear simulations performed using local parameters from various reference surfaces in STEP-EC-HD. In detail, we include several parameter scans at $\Psi_{n}=0.49$ with $\gamma_{E}\in\{0.05\,,0.1\,,0.2\}\,c_{s}/a$, $q\in\{3\,,3.5\,,4.0\,,4.5\}$, $\beta\in\{0.005\,,0.02\,,0.09\}$, and $L_{p,\mathrm{ref}}/L_{p}\in\{0.8\,,1.0\,,1.2\}$ from Giacomin et al. (2024) and an additional scan with $\hat{s}\in\{0.3\,,0.6\,,1.2\}$ for $\gamma_{E}=0$. The database also includes the simulations at $\Psi_{n}=0.36$ ($\gamma_{E}=0$ and $\gamma_{E}=0.05\,c_{s}/a$) and at $\Psi_{n}=0.8$ ($\gamma_{E}=0.1\,c_{s}/a$) described in Sec. 2.2, as well as an additional scan with $q\in\{3.0\,,3.5\,,4.0\,,4.5\}$ for $\gamma_{E}=0$ at $\Psi_{n}=0.36$.
@@ 4.1 Implementation of the reduced model in T3D
where $n_{s}$ and $p_{s}$ are the density and pressure of the species $s$, $\Psi$ is the flux label, $V$ is the volume enclosed by the surface $\Psi$, $V^{\prime}$ is the derivative of $V$ with respect to $\Psi$, $\Gamma_{s}$ and $Q_{s}$ are the total particle and heat fluxes of a given species $s$ including the classical, neoclassical and turbulent contributions, and $S_{n,s}$ and $S_{p,s}$ are the particle and the heat sources. The overline denotes a flux surface average. The term proportional to $\nu_{su}$ is the collisional energy exchange term, with $\nu_{su}$ taken from Huba (1998). Density and temperature profiles are discretised on a uniformly spaced radial grid, and gradients are computed using a sixth order finite difference scheme on interior radial grid points and a second order finite difference scheme at the boundary. A non-zero Dirichlet boundary condition is applied at the outermost radial grid point, while the physical zero flux condition is applied at the magnetic axis [^10] The magnetic axis is not included in the simulation domain. The first radial grid point is located at $r=\Delta r/2$ where $\Delta r$ is the radial grid spacing.. Eqs. (10) and (11) are evolved using a semi-implicit method. Further details on the transport equations and on their numerical implementation are reported in Barnes et al. (2010).
Evaluating the fluxes $\Gamma_{s}$ and $Q_{s}$ in Eqs. (10) and (11) requires a transport model. For instance, the coupling between T3D and GX (Mandell et al., 2022), a local gyrokinetic code developed to run natively on Graphics Processing Units (GPUs), allows T3D to submit nonlinear gyrokinetic simulations at each radial location to compute $\Gamma_{s}$ and $Q_{s}$ from first-principles simulations. A total number of $(N_{r}-1)(N_{p}+1)$ simulations per iteration is required, where $N_{r}$ is the number of radial grid points and $N_{p}$ is the number of evolved profiles. The quasi-neutrality condition is applied to impose the density value of the last evolved species. Our T3D simulations for STEP use the quasi-linear reduced transport flux equations (8) and (9), which have been implemented in T3D, and exploit a coupling to GS2 for the linear calculations required to compute the quasi-linear fluxes.
@@ 4.2.1 Transport simulation set-up and assumptions
Our T3D simulations use six radial grid points uniformly distributed over the radial domain, extending from the magnetic axis to $r/a=0.9$ close to the pedestal top. At each radial position and for each profile, a total of $n_{k_{y}}\times n_{\theta_{0}}=12\times 6=72$ linear modes are evolved, using a non-uniform grid spacing in $k_{y}$ and $\theta_{0}$ to improve resolution at low $k_{y}$ and low $\theta_{0}$, which is the region contributing most strongly to the quasi-linear metric. A minimum $k_{y}\rho_{s}$ is imposed to resolve only toroidal mode numbers $n>1$, and its value depends on radial location and plasma profiles. Each mode is evolved by an independent GS2 simulation in ballooning space. GS2 simulations are performed using 33 grid points along $\theta$, 24 pitch-angles and 10 energy grid points. The total number of GS2 simulations in ballooning space carried out at each step is $N=(N_{r}-1)(N_{p}+1)n_{k_{y}}n_{\theta_{0}}=1440$. The GS2 simulations are performed in parallel depending on the number of available MPI tasks. The auxiliary heating power density profile from electron cyclotron resonant heating ($P_{\mathrm{EC}}=150$ MW in total), radiated power density profile ($P_{\mathrm{rad}}=340$ MW in total), and particle source from pellet injection (corresponding to a fueling rate $\sim 10^{22}$ particles/s) are taken from the JINTRAC simulation of STEP-EC-HD (see Tholerus et al. (2024) for details), and kept constant throughout the T3D transport calculation. Fusion $\alpha$-particle heating and collisional energy exchange contributions to Eq. (11) evolve with the density and temperature profiles and are calculated self-consistently. The ion density is multiplied by a dilution factor, $f_{\text{He}}=0.9$, when computing the fusion power in order to account for the dilution effect from helium ash. Neoclassical transport is included in the following T3D simulations and is computed using the neoclassical code NEO (Belli & Candy, 2008).
@@ 4.2.2 Transport simulation results
Figure 14: Comparison of initial (dotted lines) and final (solid lines) temperature (a), density (c) and pressure (e) profiles as well as their inverse gradient scale length [(b), (d) and (f)]. The initial profiles are taken from the JINTRAC analysis (Tholerus et al., 2024). The round markers denotes the position of the radial grid points in the T3D grid. The value of the density and pressure at the outermost radial grid point is imposed by the finite Dirichlet boundary condition. The shaded area represents the profile variation corresponding to a $\pm$40variation in $\gamma_{E}$ and a $\pm$20variation in $Q_{0}$ and $\alpha$. The solid line represents an interpolation through the T3D radial grid points.
@@
The final T3D electron density is close to the initial profile for $r/a\leq 0.6$, but higher for $r/a>0.6$: i.e. there is a small increase in the edge density gradient for $r/a\simeq 0.8$ (which will be discussed later) and a relaxation of the gradient around $r/a\simeq 0.65$. The particle confinement time evaluated from the T3D steady state solution is $\tau_{p}=\int n_{e}\mathrm{d}V/\int S_{n}\mathrm{d}n=12.5$ s, which is close to that assumed in STEP-EC-HD (Tholerus et al., 2024).
@@
We highlight that the T3D steady state ion temperature and density gradients at $r/a\simeq 0.8$ are larger than in the initial condition from STEP-EC-HD. This may appear counter-intuitive, especially given that the reduced model heat and particle fluxes are much larger than the available sources at the STEP-EC-HD initial condition, as illustrated in Fig.16. On the other hand, the dependence of the turbulent fluxes on the temperature and density gradients is non-monotonic in STEP-EC-HD because of the stabilizing $\beta^{\prime}$ effect, as pointed out in Giacomin et al. (2024). The profile evolution in the initial phase leads to a steepening of temperature and density profile in the outer core region, which causes a local increase of the $\beta^{\prime}$ value and a subsequent reduction of turbulent fluxes, thus sustaining larger density and temperature gradients in the outer core region. We note, however, that the variation of other parameters, such as $\beta$ and collisionality, the evolution of the fusion power and the contribution from the neoclassical fluxes make the dynamics regulating the profile evolution strongly nonlinear: this effect cannot be solely explained by the dependence of turbulent fluxes on gradients.
@@
Fig. 14 shows also a comparison between the initial and final pressure profile: the reduced final pressure for $r/a<0.6$ is consistent with lower $T_{e}$ and $T_{i}$ in this region of the plasma; for $r/a>0.6$ both the electron and ion final pressure profiles are larger than the initial profiles, mostly because of the increased density. If we do not separate transport and radiation losses, the energy confinement time using the final steady state pressure profile gives $\tau_{E}=\frac{3}{2}\int p\mathrm{d}V/P_{\mathrm{net}}=1.5$ s, with $P_{\mathrm{net}}=P_{\mathrm{\alpha}}+P_{\mathrm{EC}}$. If we exclude radiation losses from the net loss power using $P^{*}_{\mathrm{net}}=P_{\mathrm{\alpha}}+P_{\mathrm{EC}}-P_{\mathrm{rad}}$, this gives $\tau^{*}_{E}=\frac{3}{2}\int p\mathrm{d}V/P^{*}_{\mathrm{net}}=4.6$ s. These energy confinement times are similar to those assumed in the original STEP flat-top operating point [^12] The T3D steady state energy confinement time is slightly higher than in the original STEP-EC-HD ($\tau_{E}=1.24$ s and $\tau_{E}^{*}=4.20$ s) because of the larger density in the outer core, which leads to a slightly larger $\int p\mathrm{d}V$. (Tholerus et al., 2024).
@@
The shaded area in Fig. 14 indicates the range of profiles variation obtained from T3D simulations with local values of $\gamma_{E}$ varied by $\pm$40from the nominal value, and also includes additional T3D simulations where $Q_{0}$ and $\alpha$ are varied by $\pm$20to assess the impacts of uncertainties in the reduced model parameters. The variation in the density and temperature profiles is comparatively small, which may seem surprising given the large sensitivity of fluxes to $\gamma_{E}$ in STEP-EC-HD. This relatively weak sensitivity of the kinetic profiles on the shearing rate can be explained (i) by T3D converging to a regime where the linear hybrid-KBM instability is mitigated, especially at low $k_{y}$; and (ii) by large transport stiffness, where small variations in the temperature and density gradients lead to large changes in the turbulent flux. High transport stiffness improves the robustness of the T3D steady state.
@@
Figure 17: Fusion power evolution in the reference T3D simulation. The fusion power is normalized to its initial value of 1.7 GW.
The evolution of the fusion power is shown in Fig. 17. Consistently with the pressure profile evolution, the fusion power rapidly decreases in the initial phase, which is followed by a slower evolution. The final fusion power is approximately 10% smaller than its initial value.

### v2

@@ 1 Introduction
High-fidelity tools to evolve plasma equilibrium profiles on the transport time scale have been developed in recent years by coupling transport solvers to local or global gyrokinetic codes. These codes, such as TANGO-GENE (Di Siena et al., 2022), PORTALS-CGYRO (Rodriguez-Fernandez et al., 2022) and TRINITY (Barnes et al., 2010), exploit the enormous separation between the timescales associated with turbulence and equilibrium transport, to evolve plasma profiles self-consistently with the turbulence. For instance, PORTALS-CGYRO has been recently applied to perform the first flux-matched nonlinear gyrokinetic predictions of SPARC profiles, where Gaussian process regression techniques have been applied to optimize the transport solver (Rodriguez-Fernandez et al., 2022). Naturally high-fidelity models require considerably more computational resources than using reduced transport models [^1] This is true even when Gaussian process regression techniques are used to reduce the number of solver iterations (Rodriguez-Fernandez et al., 2022).; and their application to STEP is computationally prohibitively expensive, because each of the many nonlinear local simulations required by a transport solver costs more than $10^{5}$ core-hours. Furthermore, starting from an initial state that strongly violates power balance increases significantly the number of nonlinear gyrokinetic simulations required for profile convergence (power balance is very poorly satisfied at several radial locations in STEP-EC-HD as assumed in scenario design with JINTRAC, which provides the initial condition in our flux-driven transport calculations). There is therefore an increasing interest in developing low-cost and high-fidelity transport solvers that can be exploited in integrated modeling, such as the fast and accurate turbulent transport model recently developed by Citrin et al. (2023), where the saturation rule, calibrated to a set of nonlinear GENE simulations, is combined with a neural network regression to perform flux-matched predictions of the ITER baseline scenario.
@@ 2.1 Linear simulation results
Figure 1: Growth rate (a) and mode frequency (b) as functions of the binormal wave vector $k_{y}$ at various radial surfaces of STEP-EC-HD. The inset in (b) shows the transition to positive frequency values at low $k_{y}$ of $\Psi_{n}=0.15$.
@@
On the innermost surface at $\Psi_{n}=0.15$, however, the dominant modes have negative frequency and very different parallel mode structures. These are illustrated in Fig. 3 [(a) and (b)] for the fastest growing mode at $k_{y}\rho_{s}\simeq 0.84$: $\delta\phi$ has odd parity and is highly extended in $\theta$ whilst $\delta A_{\parallel}$ is very localised at $\theta=0$ with even parity. Amplitudes of $e\delta\phi/T_{e}$ and $\delta A_{\parallel}/(\rho_{s}B_{0})$ are comparable, indicating that the mode is strongly electromagnetic. These negative frequency modes propagate in the electron diamagnetic drift direction, and have the typical properties expected of MTMs (Applegate et al., 2004). On the same surface at the lowest $k_{y}$, on the other hand, the dominant mode switches to a positive frequency (see the inset in Fig. 1 (b)) and very different parallel mode structures in $\delta\phi$ and $\delta A_{\parallel}$. The parallel mode structure of a low $k_{y}$ mode with positive frequency is shown in Fig. 3 [(c) and (d)]: the parallel structure closely resemble Fig. 2, indicating that the hybrid-KBM is linearly dominant at low $k_{y}$. Although these low $k_{y}$ hybrid-KBMs are only weakly unstable, they nevertheless contribute significantly to the turbulent transport, as will be discussed in Sec. 2.2.
@@ 3.1 Quasi-linear reduced transport model
Species heat and particle fluxes are obtained (from Eqs. (8) and (9) below) using the total model heat flux, $Q_{\mathrm{ql}}=Q_{0}\Lambda^{\alpha}$ defined in Eq. (2) that sets the turbulence saturation amplitude, together with the quasi-linear weights $Q_{\mathrm{l},s}/Q_{\mathrm{l}}$ and $\Gamma_{\mathrm{l},s}/Q_{\mathrm{l}}$, where $Q_{\mathrm{l},s}$ and $\Gamma_{\mathrm{l},s}$ are the linear heat and particle fluxes for species $s$, and $Q_{\mathrm{l}}=\sum_{s}Q_{\mathrm{l},s}$ is the total linear heat flux; all linear fluxes are taken from the final time point in a converged linear simulation for a given mode (i.e. $k_{y},\theta_{0}$). The quasi-linear weights apportion contributions to transport fluxes for each species for a given mode (depending on $k_{y}$ and $\theta_{0}$) must be included in the integrands of Eqs (8) and (9) to obtain the total species heat and particle fluxes:
@@ 3.2 Application to STEP relevant regimes
The parameters $Q_{0}$ and $\alpha$, which link $\Lambda$ to the quasi-linear heat flux $Q_{\mathrm{ql}}$ via Eq. (2), are tuned here to optimise the description of the total heat transport from hybrid-KBM turbulence in STEP. Suitable values are determined by fitting the nonlinear saturated heat fluxes from a set of nonlinear gyrokinetic simulations for STEP relevant local equilibria. This database contains nonlinear simulations performed using local parameters from various reference surfaces in STEP-EC-HD. The parameter scans reported in Kennedy et al. (2023) and Giacomin et al. (2024) already cover many of the most important local parameters, including pressure gradient, $\beta$, flow shear, $q$ and $\hat{s}$, to which hybrid-KBM turbulence is sensitive. Those scans provide the bulk of the database used to develop the quasi-linear model. In detail, we include several parameter scans at $\Psi_{n}=0.49$ with $\gamma_{E}\in\{0.05\,,0.1\,,0.2\}\,c_{s}/a$, $q\in\{3\,,3.5\,,4.0\,,4.5\}$, $\beta\in\{0.005\,,0.02\,,0.09\}$, and $L_{p,\mathrm{ref}}/L_{p}\in\{0.8\,,1.0\,,1.2\}$ from Giacomin et al. (2024) and an additional scan with $\hat{s}\in\{0.3\,,0.6\,,1.2\}$ for $\gamma_{E}=0$. The database also includes the simulations at $\Psi_{n}=0.36$ ($\gamma_{E}=0$ and $\gamma_{E}=0.05\,c_{s}/a$) and at $\Psi_{n}=0.8$ ($\gamma_{E}=0.1\,c_{s}/a$) described in Sec. 2.2, as well as an additional scan with $q\in\{3.0\,,3.5\,,4.0\,,4.5\}$ for $\gamma_{E}=0$ at $\Psi_{n}=0.36$.
@@ 4.1 Implementation of the reduced model in T3D
where $n_{s}$ and $p_{s}$ are the density and pressure of the species $s$, $\Psi$ is the flux label, $V$ is the volume enclosed by the surface $\Psi$, $V^{\prime}$ is the derivative of $V$ with respect to $\Psi$, $\Gamma_{s}$ and $Q_{s}$ are the total particle and heat fluxes of a given species $s$ including the classical, neoclassical and turbulent contributions, and $S_{n,s}$ and $S_{p,s}$ are the particle and the heat sources. The overline denotes a flux surface average. The term proportional to $\nu_{su}$ is the collisional energy exchange term, with $\nu_{su}$ taken from Huba (1998). Density and temperature profiles are discretised on a uniformly spaced radial grid, and gradients are computed using a sixth order finite difference scheme on interior radial grid points and a second order finite difference scheme at the boundary. A non-zero Dirichlet boundary condition is applied at the outermost radial grid point, while the physical zero flux condition is applied at the magnetic axis [^10] The first radial grid point is located at $r=\Delta r/2$ where $\Delta r$ is the radial grid spacing. Fluxes and gradients are evaluated on interior radial grid points (apart from the magnetic axis where fluxes vanish) up to the outermost radial grid point $r_{N_{r}}-\Delta r/2$. The value of density and temperature at the magnetic axis can be evaluated using the zero flux on-axis boundary condition.. Eqs. (10) and (11) are evolved using a semi-implicit method. Further details on the transport equations and on their numerical implementation are reported in Barnes et al. (2010).
Evaluating the fluxes $\Gamma_{s}$ and $Q_{s}$ in Eqs. (10) and (11) requires a transport model. For instance, the coupling between T3D and GX (Mandell et al., 2024), a local gyrokinetic code developed to run natively on Graphics Processing Units (GPUs), allows T3D to submit nonlinear gyrokinetic simulations at each radial location to compute $\Gamma_{s}$ and $Q_{s}$ from first-principles simulations. A total number of $(N_{r}-1)(N_{p}+1)$ simulations per iteration is required, where $N_{r}$ is the number of radial grid points and $N_{p}$ is the number of evolved profiles. We note that $N_{p}+1$ evaluations are required at each radial location to compute the local flux derivatives with respect to each evolved profile gradient (see Barnes et al. (2010) for details). The quasi-neutrality condition is applied to impose the density value of the last evolved species. Our T3D simulations for STEP use the quasi-linear reduced transport flux equations (8) and (9), which have been implemented in T3D, and exploit a coupling to GS2 for the linear calculations required to compute the quasi-linear fluxes.
@@ 4.2.1 Transport simulation set-up and assumptions
Our T3D simulations use six radial grid points uniformly distributed over the radial domain, extending from the magnetic axis to $r/a=0.9$ close to the pedestal top. At each radial position and for each profile, a total of $n_{k_{y}}\times n_{\theta_{0}}=12\times 6=72$ linear modes are evolved, using a non-uniform grid spacing in $k_{y}$ and $\theta_{0}$ to improve resolution at low $k_{y}$ and low $\theta_{0}$, which is the region contributing most strongly to the quasi-linear metric. A minimum $k_{y}\rho_{s}$ is imposed to resolve only toroidal mode numbers $n>1$, and its value depends on radial location and plasma profiles. Each mode is evolved by an independent GS2 simulation in ballooning space. GS2 simulations are performed using 33 grid points along $\theta$, 24 pitch-angles and 10 energy grid points. The total number of GS2 simulations in ballooning space carried out at each step is $N=(N_{r}-1)(N_{p}+1)n_{k_{y}}n_{\theta_{0}}=1440$, where $N_{p}=3$ as $n$, $T_{e}$ and $T_{i}$ are evolved here. The GS2 simulations are performed in parallel depending on the number of available MPI tasks. The auxiliary heating power density profile from electron cyclotron resonant heating ($P_{\mathrm{EC}}=150$ MW in total), radiated power density profile ($P_{\mathrm{rad}}=340$ MW in total), and particle source from pellet injection (corresponding to a fueling rate $\sim 10^{22}$ particles/s) are taken from the JINTRAC simulation of STEP-EC-HD (see Tholerus et al. (2024) for details), and kept constant throughout the T3D transport calculation. Fusion $\alpha$-particle heating and collisional energy exchange contributions to Eq. (11) evolve with the density and temperature profiles and are calculated self-consistently. The ion density is multiplied by a dilution factor, $f_{\text{He}}=0.9$, when computing the fusion power in order to account for the dilution effect from helium ash. Motivated by the presence of subdominant MTM turbulence in the STEP initial condition (Giacomin et al., 2024), a small additional Rechester-Rosenbluth-like contribution is included in the electron heat flux proportional to $(a/L_{T_{e}})^{3}$ (see Doerk et al. (2011)). This helps convergence, while it has a negligible impact in the final T3D steady state. Neoclassical transport is included in the following T3D simulations and is computed using the neoclassical code NEO (Belli & Candy, 2008).
@@ 4.2.2 Transport simulation results
Figure 14: Comparison of initial (dotted lines) and final (solid lines) temperature (a), density (c) and pressure (e) profiles as well as their inverse gradient scale length [(b), (d) and (f)]. The initial profiles are taken from the JINTRAC analysis (Tholerus et al., 2024). The round markers denotes the position of the radial grid points in the T3D grid. The value of the density and pressure at the outermost radial grid point is imposed by the finite Dirichlet boundary condition. The shaded area represents the profile variation corresponding to a $\pm$40variation in $\gamma_{E}$ and a $\pm$20variation in $Q_{0}$ and $\alpha$. The solid line represents an interpolation through the T3D radial grid points. The particle source (star markers) as well as the electron (square markers) and ion (diamond markers) total power sources are shown in (c) and (e).
@@
The final T3D electron density is close to the initial profile for $r/a\leq 0.6$, but higher for $r/a>0.6$: i.e. there is a small increase in the edge density gradient for $r/a\simeq 0.8$ and a relaxation of the gradient around $r/a\simeq 0.65$. We note that the density increase occurs in proximity of the maximum density source due to pellet injection, shown in Fig. 14 (c). The particle confinement time evaluated from the T3D steady state solution is $\tau_{p}=\int n_{e}\mathrm{d}V/\int S_{n}\mathrm{d}n=12.5$ s, which is close to that assumed in STEP-EC-HD (Tholerus et al., 2024). We highlight that the T3D steady state ion temperature and density gradients at $r/a\simeq 0.8$ are larger than in the initial condition from STEP-EC-HD. This may appear counter-intuitive, especially given that the reduced model heat and particle fluxes are much larger than the available sources at the STEP-EC-HD initial condition, as illustrated in Fig.16. On the other hand, the dependence of the turbulent fluxes on the temperature and density gradients is non-monotonic in STEP-EC-HD because of the stabilizing $\beta^{\prime}$ effect, as pointed out in Giacomin et al. (2024). The profile evolution in the initial phase leads to a steepening of temperature and density profile in the outer core region, which causes a local increase of the $\beta^{\prime}$ value and a subsequent reduction of turbulent fluxes, thus sustaining larger density and temperature gradients in the outer core region. We note, however, that the variation of other parameters, such as $\beta$ and collisionality, the evolution of the fusion power and the contribution from the neoclassical fluxes make the dynamics regulating the profile evolution strongly nonlinear: this effect cannot be solely explained by the dependence of turbulent fluxes on gradients.
@@
Fig. 14 shows also a comparison between the initial and final pressure profile: the reduced final pressure for $r/a<0.6$ is consistent with lower $T_{e}$ and $T_{i}$ in this region of the plasma; for $r/a>0.6$ both the electron and ion final pressure profiles are larger than the initial profiles, mostly because of the increased density. If we do not separate transport and radiation losses, the energy confinement time using the final steady state pressure profile gives $\tau_{E}=\frac{3}{2}\int p\mathrm{d}V/P_{\mathrm{net}}=1.5$ s, with $P_{\mathrm{net}}=P_{\mathrm{\alpha}}+P_{\mathrm{EC}}$. If we exclude radiation losses from the net loss power using $P^{*}_{\mathrm{net}}=P_{\mathrm{\alpha}}+P_{\mathrm{EC}}-P_{\mathrm{rad}}$, this gives $\tau^{*}_{E}=\frac{3}{2}\int p\mathrm{d}V/P^{*}_{\mathrm{net}}=4.6$ s. These energy confinement times are similar to those assumed in the original STEP flat-top operating point [^12] The T3D steady state energy confinement time is slightly higher than in the original STEP-EC-HD ($\tau_{E}=1.24$ s and $\tau_{E}^{*}=4.20$ s) because of the larger density in the outer core, which leads to a slightly larger $\int p\mathrm{d}V$. (Tholerus et al., 2024). We highlight, however, that the radial domain evolved by these T3D simulations does not include the pedestal region. The outermost radial grid point boundary condition of the density and temperature is taken from near the pedestal top of the JINTRAC initial profiles (see Fig. 14); the pedestal model used in the JINTRAC modelling of this STEP flat-top operating point is described in Tholerus et al. (2024). Any variation of the pedestal top, and therefore of the outer boundary condition, is likely to impact core transport and the energy confinement time. Assessing the sensitivity on the outer boundary condition, i.e. on the density and temperature values at the pedestal top, is clearly a very important topic for future investigation.
@@
The shaded area in Fig. 14 indicates the range of profiles variation obtained from T3D simulations with local values of $\gamma_{E}$ varied by $\pm$40from the nominal value, and also includes additional T3D simulations where $Q_{0}$ and $\alpha$ are varied by $\pm$20to assess the impacts of uncertainties in the reduced model parameters. The variation in the density and temperature profiles is comparatively small, which may seem surprising given the large sensitivity of fluxes to $\gamma_{E}$ in STEP-EC-HD. This relatively weak sensitivity of the kinetic profiles on the shearing rate can be explained (i) by T3D converging to a regime where the hybrid-KBM turbulence is closer to marginal, especially at low $k_{y}$, in the transport steady state, owing reductions of the density gradient in the core and increases in $\beta^{\prime}$ at the edge; and (ii) by large transport stiffness, where small variations in the temperature and density gradients lead to large changes in the turbulent flux. High transport stiffness improves the robustness of the T3D steady state.
@@
Figure 17: (a) Power source radial profile in the final T3D steady state of the auxiliary heating ($S_{\mathrm{aux}}$), $\alpha$-particle heating ($S_{\alpha,e}$ and $S_{\alpha,i}$), collisional exchange ($S_{\mathrm{coll},e}=-S_{\mathrm{coll},i}$) and radiation ($S_{\mathrm{rad}}$). (b) Fusion power evolution in the reference T3D simulation. The fusion power is normalized to its initial value of 1.7 GW.
The heating power density radial profiles from auxiliary and $\alpha$-particle heating, collisional exchange, and radiation are shown in Fig. 17 (a) for the final T3D steady state, while the evolution of the fusion power is shown in Fig. 17 (b). Consistently with the pressure profile evolution, the fusion power rapidly decreases in the initial phase, which is followed by a slower evolution. The final fusion power is approximately 10% smaller than its initial value.